1. Compute the molecular weight of a single, (1), polystyrene (PS) repeat unit.

2. Compute the degree of polymerization for a polystyrene (PS) polymer with a number-average molecular weight of 279,799 g/mol.

3. What is the number-average molecular weight of a PC polymer that features a degree of polymerization of 1,890 ?

4. An alternating poly(ethylene-propylene) copolymer has a molecular weight of 294,419 g/mol. How many ethylene repeat units are in one average length chain of this polymer ?

5. Compute the end-to-end separation (direct distance between molecule ends), in Angstroms, of a coiled PVC molecule with a number-average molecular weight of 215,891 g/mol, given that the tetrahedral bond angle is 109.5o and the carbon-carbon bond length is 1.54 Angstroms.

6. Compute the length, in microns, of a fully extended PVC molecule with a molecular weight of 189,442 g/mol, given that the tetrahedral bond angle is 109.5o and the carbon-carbon bond length is 1.54 Angstroms.
Answer #1

SOLN 1

Structure of polystyrene is

According to Chegg Policy only 1 question per post is allowed. please post the rest question seperately. However, I am solving 2nd and 3rd also

SOLN 2

DP = Mn / Mw

where Mn = number average molar mass

DP = 279,799 / 104.14 = 2686.7

SOLN 3

if DP = 1890

then 1890 = Mn / 104.14

hence Mn = 196824.6 g/mol

PLEASE LEAVE THE POSITIVE FEEDBACK IF YOU LIKE THE SOLUTION